OpenStudy (itrymath):

Create your own equation written in standard form. Determine any point that is a solution and will be on this line. Justify your answer algebraically. Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs. Rewrite the equation using function notation. Graph the function and show the intercepts.

OpenStudy (itrymath):

@CamPayne

OpenStudy (campayne):

First step: Write an equation for a line such as... y = 2x + 4 This pick a point, like (1, 4) Plug it into equation (4) = 2(1) + 4 4 = 2 + 4 4 = 6 Not correct, so (1,4) isn't a point on the line See?

OpenStudy (danjs):

standard form is -- ax + by = c

OpenStudy (danjs):

pick values for a,b,and c...

OpenStudy (itrymath):

DanJs guide me or someone guide me

OpenStudy (itrymath):

okay

OpenStudy (itrymath):

2x+3y=10

OpenStudy (danjs):

ok, that is the first part.. for the second part, they want any point (x,y) that is on that line... 2x + 3y = 10 you can pick any x value, then solve for what y would be, to get (x,y)

OpenStudy (itrymath):

um i choose you

OpenStudy (itrymath):

jk x=2

OpenStudy (danjs):

ok, put the 2 in for x, and solve for y

OpenStudy (itrymath):

4+3y=10

OpenStudy (danjs):

right, x=2, not solve for y=

OpenStudy (itrymath):

y= 10/3 - 4/3

OpenStudy (danjs):

right, that comes out to 6/3 or y=2 It is better to move the whole numbers around first before dividing if you want. 4 + 3x = 10 3x = 6 x=2

OpenStudy (itrymath):

DOPE! i thought i was wrong for a second because i was following the steps to get f(x)

OpenStudy (danjs):

it still works, just more messy.., so you choose x=2, and calculated y=2, so you get a point on the line (x,y)=(2,2)

OpenStudy (itrymath):

yup

OpenStudy (danjs):

The x and y intercepts are the points where the line crosses the x-axis, and y=axis. For any point on the x-axis, the y value is 0. you see that>

OpenStudy (itrymath):

DanJS you're following the questions it is asking right?

OpenStudy (itrymath):

Just making sure

OpenStudy (danjs):

yeah, the first one, standard form ax+by=c 2x + 3y = 10 the second, a point (x,y) on the line, you choose x=2, and got y=2, (x,y) = (2,2) the third, find the x and y axis intercept points....

OpenStudy (itrymath):

Explain how to find the x- and y- intercept(s) using complete sentences. Write the intercepts as ordered pairs.

OpenStudy (danjs):

ok, you see how any point on the x-axis line has a y value zero... and any point on the y-axis has an x value x=0?? 

OpenStudy (itrymath):

yes

OpenStudy (danjs):

The x and y intercepts for your line, are where it crosses each axis.

OpenStudy (itrymath):

yes

OpenStudy (danjs):

to get where it crosses the x-axis, you set the y value to zero...and to get the y-axis intercept, you set the x value to zero....the two points are y-intercept ---- (0 , y) x-intercept -----(x , 0) you see that?

OpenStudy (danjs):

i am here

OpenStudy (danjs):

So you do the same thing like you did finding a point(x,y), except this time, you want when x=0 and when y=0

OpenStudy (danjs):

2x+3y=10

OpenStudy (itrymath):

and what will that give me? that would give me 0=10...

OpenStudy (danjs):

When x=0, you get the y-axis intercept... 2x + 3y = 10 2*0 + 3y = 10 3y = 10 y = 10/3 When x=0, y=10/3 , y-intercept (x,y)=(0, 10/3 )

OpenStudy (danjs):

Same for the x-intercept, except this time put y=0

OpenStudy (itrymath):

ooo yeah sorry i plugged in 0 in both palces at the same time

OpenStudy (itrymath):

x=5

OpenStudy (itrymath):

y=5**

OpenStudy (danjs):

you had it the first time, if y=0, x=5

OpenStudy (danjs):

So you have the x and y intercepts (5,0) and (0 , 10/3)

OpenStudy (itrymath):

okay

OpenStudy (itrymath):

it says reqwrite the function using notation .... and graph the function

OpenStudy (itrymath):

reqrite

OpenStudy (danjs):

2x + 3y = 10 For a function of y in terms of x, y=f(x), solve your equation for y

OpenStudy (danjs):

then you have a function y = f(x), where you pick an x value, and it gives you the y value

OpenStudy (itrymath):

explain please

OpenStudy (itrymath):

so i put in any numbers or the numbers we are using

OpenStudy (danjs):

You take your equation that is in standard form ax + by = c, 2x + 3y = 10 and solve that for y

OpenStudy (danjs):

2x + 3y = 10 3y = -2x + 10 \[\large y=\frac{ -2 }{ 3 }x+\frac{ 10 }{ 3 }\]

OpenStudy (itrymath):

the points were (2,2) so i plug those in

OpenStudy (itrymath):

oh okay

OpenStudy (itrymath):

got it

OpenStudy (danjs):

as a function y in terms of x \[y=f(x)\] \[f(x)=\frac{ -2 }{ 3 }x + \frac{ 10 }{ 3 }\]

OpenStudy (itrymath):

how do i graph that?

OpenStudy (danjs):

you only need 2 points to graph a line, i would use the two axis intercepts you found

OpenStudy (itrymath):